OpenStudy (anonymous):

what is the derivative of: y=sin^-1 x , x = 3/5

OpenStudy (anonymous):

\[-\sin ^{-2}x*\cos ^{-1}x\] that's the derivative but you should use a calculator to find the answer.

OpenStudy (anonymous):

You know what the derivative of arcsine is? :)

OpenStudy (anonymous):

That's right :D \[\huge \frac{d}{dx}\sin^{-1}x=\frac{1}{\sqrt{1-x^2}}\] Now just plug in x = 3/5 and you'll be all set :D

OpenStudy (anonymous):

Awesome :)

OpenStudy (anonymous):

Yes, pretty much. But I think you won't get as nice an answer. Oh well, that's life, you get the nice things and the not-so-nice things :D

OpenStudy (anonymous):

first, we might want to get the derivative of arccos(4x) and I hope you got \[\huge \frac{d}{dx}\cos^{-1}4x=\frac{-4}{\sqrt{1-x^2}}\]

OpenStudy (anonymous):

Now just plug in x = 1/5 :D

OpenStudy (anonymous):

\[-4\div2\sqrt{6\div5}\]

OpenStudy (anonymous):

Let's do it from the top, instead, because I'm confused ;) \[\frac{-4}{\sqrt{1-\left(\frac{1}{5}\right)^2}}\]

OpenStudy (anonymous):

\[\large \frac{-4}{\sqrt{\frac{25}{25}-\frac{1}{25}}}\]

OpenStudy (anonymous):

\[\huge \frac{-4}{\sqrt{\frac{24}{25}}}=\frac{-4\sqrt{25}}{\sqrt{24}}\]

OpenStudy (anonymous):

Well, you might consider... \[\huge \frac{-4(5)}{2\sqrt{6}}\] and do the cancel stuffs :D

OpenStudy (anonymous):

You got it? Awesome :D

OpenStudy (anonymous):

:)